Adult tickets for a play cost $14 and child tickets cost $4. If there were 28 people at a performance and the theater collected $242 from ticket sales, how many children attended the play? Solve this problem using algebraic equations.
Thank you.

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

let x=children, so 28-x=adults ... 4(x)+14(28-x)=242

Answer by stanbon(75887)

Adult tickets for a play cost $14 and child tickets cost $4. If there were 28 people at a performance and the theater collected $242 from ticket sales, how many children attended the play? Solve this problem using algebraic equations.
---------------------
Let # of adult tickets be "x" ; Value of these is 14x dollars
---------------------
# of child tickets is "28-x" ; value of these is 4(28-x) = 112-4x dollars
-----------------------
EQUATION:
value + value = 242 dollars
14x + 112-4x = 242
10x = 130
x = 13 (# of adult tickets sold)
28-x = 15 (# of child tickets sold)
===============
Cheers,
Stan H.